# Perspective on Moreau–Yosida Regularization in Density-Functional Theory

### v1

@@ II.4 Formal density-functional theory
But it is also known since a long time that $v$-representability, even if it is secured for a dense set, is a problematic notion since many densities $\rho\in D$ will not have this property [33; 34]. Even worse, in the standard setting of DFT with density space $L^{1}\cap L^{3}$, arbitrarily close to any $v$-representable density there lies a non-$v$-representable one, which makes $F^{\lambda}(\rho)$ discontinuous and non-differentiable [35]. In many cases (including standard DFT) the set of $v$-representable densities is known only implicitly (exceptions are lattice systems [36; 32; 37] and particles on a one-dimensional torus [17; 18]). We will later see how MY regularization offers a way out of this dilemma.
@@
The Hohenberg–Kohn theorem [28] states that, for a $v$-representable density $\rho\in D$, the corresponding potential is unique up to addition of constants. In standard DFT, a rigorous proof of the Hohenberg–Kohn theorem has been achieved for $v\in L^{p}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$, with $p>2$ [38] and it extends to closely related settings [39]. This leaves $3/2\leq p\leq 2$ as undecided (the lower bound is given by the argument for self-adjoint Hamiltonians as noted before), although results on unique continuation [40; 41] suggest that the Hohenberg–Kohn theorem holds also in this range. In non-standard DFT settings, like finite lattices and when currents are included, counterexamples are known [32; 42].
@@
Combining the assumptions from this section and from Section II.1 (which will later serve as the basis for MY regularized DFT), we arrive at the following important result:
@@
Proposition. Assumption 1 together with Assumption 4 with a potential space $X^{*}\subset L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$ imply that every $v\in X^{*}$ supports a ground state—that is, for all $v\in X^{*}$, the superdifferential $\partial E^{\lambda}(v)$ is non-empty.
@@
This claim can be seen as follows. The choice $L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$ as a potential space makes $E^{\lambda}(v)$ continuous [3, Th. 3.1(iii)], implying that each continuously embedded $X^{*}\subset L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$ gives a continuous energy functional $E^{\lambda}(v)$. But a continuous and concave functional on $X^{*}$ admits a non-empty subdifferential belonging to $X^{\ast\ast}$ everywhere in the interior of its domain (here all $X^{\ast}$) [5, Prop. 2.36]. In general, it only holds $X^{**}\supseteq X$, but with Assumption 1 we have reflexivity and thus $X=X^{**}$. By the additional requirement that $F_{\mathrm{DM}}^{\lambda}(\rho)$ is lower semicontinuous, we have $F^{\lambda}(\rho)=F_{\mathrm{DM}}^{\lambda}(\rho)$. Each $\rho\in\partial E^{\lambda}(v)$ is now in $D\subset X$ and through $v\in-\partial F^{\lambda}(\rho)=-\partial F_{\mathrm{DM}}^{\lambda}(\rho)$ it is connected to an actual ground-state solution of the Schrödinger equation as argued above. This demonstrates the claim of the proposition.
@@
We illustrate this subtle and important point with an example. Take the Hamiltonian from Eq. (23) with just a single particle and set $v=0$. Then there is no bound state and thus $\partial E^{\lambda}(v)$ should be empty (note that the $\lambda$ parameter is not needed for single particles since there is no interaction anyway). Indeed, this example is incompatible with reflexivity and lower semicontinuity assumed together. If we give up on reflexivity of the spaces, then we can find an $x\in\partial E^{\lambda}(0)$ where $x\in X^{**}\setminus X$, so it does not represent a ground state. Next, we keep reflexivity but give up on lower semicontinuity of $F_{\mathrm{DM}}(\rho)$. We pick a minimizing sequence $\rho_{i}\to x$ in Eq. (30) consisting, for example, normalized Gaussian-shaped densities that get broader with increasing $i$ and are all in $D$, while $x\neq D$. Note that this means $x=0$, which cannot hold if $X$ includes the $L^{1}$ norm (and which would make it non-reflexive), since all $\rho_{i}$ were assumed normalized to the same number and clearly $N=\|\rho_{i}\|_{L^{1}}\to 0$ is a contradiction. But it is not necessarily a contradiction in a reflexive $X$ without the $L^{1}$ norm. Now the chosen sequence fulfills $\lim F_{\mathrm{DM}}(\rho_{i})=0$ but at the same time $F_{\mathrm{DM}}(x)=+\infty$, since $x\notin D$, so lower semicontinuity cannot hold.
@@
This section summarized the most important features of DFT in the language of functional and convex analysis. The reader may wonder if Assumption 1 (or one of the two stronger ones) and Assumption 4 can be satisfied simultaneously. After all, the classical setting of DFT, with $X=L^{1}(\mathbb{R}^{3})\cap L^{3}(\mathbb{R}^{3})$ for densities and its dual $X^{\ast}=L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$ for potentials, allows for Assumption 4 but breaks reflexivity, a property not needed for a consistent formulation of DFT but necessary for our later results related to Moreau–Yosida regularization. One solution is to choose a bounded domain $\Omega$ which implies a continuous embedding $L^{p}(\Omega)\subseteq L^{q}(\Omega)$, $q<p$. Given that the important result of lower semicontinuity due to Lieb [3, Th. 4.4 and Cor. 4.5] holds with respect to the $L^{1}$ topology, it then extends to each $L^{p}(\Omega)$, $p>1$. We can then choose any $X=L^{p}(\Omega)$ with $1<p<\infty$ with dual $X^{*}=L^{p^{*}}(\Omega)$, $1/p+1/p^{*}=1$, and thus have Assumptions 2 and 4 simultaneously satisfied. By selecting $p=2$, we satisfy both Assumptions 3 and 4. The previous works of Eschrig [13] and Kvaal et al. [1] used this route, while many other works on the topic had to give up on at least one of the assumptions. Rather than switching to a bounded domain, we can use the offset potential $v_{\mathrm{o}}$ to introduce a confining potential for the whole system and thus ensure that every $v\in X^{*}$ yields a ground-state density. This, at the same time, serves to break the full translational and rotational symmetry of the setting, which is necessary to allow for any localized ground state in the first place [43].
@@ II.5 Kohn–Sham method
DFT reformulates the ground-state problem of quantum mechanics by replacing the many-particle wave function as the central variable with the one-particle density, thereby significantly reducing the complexity of the description. Yet, the usefulness of such a reduced description entirely depends on whether $F^{\lambda}(\rho)$ can be evaluated accurately and efficiently. The $F^{0}(\rho)$ of a non-interacting system amounts to just the kinetic energy and its conjugate functional $E^{0}(v)$ can be computed effectively by means of orbitals. For the interacting case, however, $F^{1}(\rho)$ is in principle unavailable since its evaluation corresponds to a quantum version of an NP-hard problem [44]. The difference $F^{1}(\rho)-F^{0}(\rho)$, which relates to the exchange–correlation functional, has moreover turned out to be difficult to model in terms of $\rho$ [45]. Thus, although the exact DFT formulation is very appealing by avoiding the full description by a wave function, further steps are needed to attain a useful reformulation. In the KS method [46], an auxiliary system of non-interacting electrons is introduced and its external potential $v_{\mathrm{s}}$ is chosen in such a way that it reproduces the same density as the interacting system. Yet, we will argue below that it is by no means clear that this approach is always possible.
The exchange–correlation potential is defined as the potential that must be added to the external potential and the Hartree potential of the auxiliary, non-interacting system in order to get the same ground-state density as the interacting system. This means that the method requires $\rho\in D$ to be simultaneously interacting and non-interacting $v$-representable, which is equivalent to $\partial F^{1}(\rho)$ and $\partial F^{0}(\rho)$ both non-empty. We already noted on the general issues regarding $v$-representability, and this additional requirement of simultaneous $v$-representability comes as a further restriction [47]. Assuming for now that this is ensured, we can take $v\in-\partial F^{1}(\rho)$ and $v_{\mathrm{s}}\in-\partial F^{0}(\rho)$ and define the exchange–correlation potential
@@
that itself is the functional derivative of the Hartree mean-field energy
@@
$v_{\mathrm{o}}+v_{\mathrm{s}}=v_{\mathrm{o}}+v+v_{\mathrm{H}}(\rho)+v_{\mathrm{xc}}(\rho),$ (43)
@@
The second step corresponds to solving the non-interacting Schrödinger equation with an external potential $v_{i+1}$ and determining a ground-state density $\rho_{i+1}$. Since the Hamiltonian is bounded below for $v_{i+1}\in X^{*}$, this approach is always possible. A more detailed convex-analytical argument for the non-emptiness of the superdifferential can be found elsewhere [14, proof of Th. 12]. The two steps of Eqs. (44) and (45) can be combined into a single step, either for the potentials or, as we will show here, for the densities,
@@
For initialization, we choose an initial density $\rho_{1}$ or, equivalently, a potential $v_{1}$ for the auxiliary system that roughly leads to the desired ground-state density $\rho$. Since the auxiliary system does not include any interactions between the particles, we (in the simplest case) choose its ground state as a Slater determinant and solve Eq. (45) by filling up the orbitals of a one-particle Schrödinger equation with external potential $v_{\mathrm{o}}+v_{i+1}$. If the procedure converges, we get
@@
so indeed both systems share the same ground-state density if in the auxiliary system an external potential $v_{\mathrm{o}}+v_{\mathrm{s}}$ is applied. Finally, in order to conveniently retrieve the interacting energy $E^{1}(v)$ from a converged density and the expressions defined above, we can just use $E^{0}(v_{\mathrm{s}})=F^{0}(\rho)+\langle v_{\mathrm{s}},\rho\rangle$ and calculate
@@
where we have introduced the exchange–correlation functional $E_{\mathrm{xc}}(\rho)=F^{1}(\rho)-F^{0}(\rho)-E_{\rm H}(\rho)$. We remark that practical KS-DFT methods are usually based on approximating $E_{\mathrm{xc}}(\rho)$. Assuming differentiability, we then have $v_{\mathrm{xc}}=\nabla E_{\mathrm{xc}}(\rho)$.
Note that this formulation of the KS method does not come with any guarantee of convergence. Indeed, in practice, an additional damping step [48; 49; 50] like taking $\rho_{i}+\alpha(\rho_{i+1}-\rho_{i})$, $\alpha<1$, as the next density is introduced to facilitate convergence. A mathematical proof of convergence has only been achieved in the MY-regularized setting, which is in any case needed to avoid the $v$-representability problem, as will be discussed in Section V.3.
@@ III Moreau–Yosida regularization in density-functional theory
Having the stage set, we now show three different ways to introduce MY regularization into DFT, which are equivalent and thus have the same outcome—namely, replacing the universal density functional $F^{\lambda}(\rho)$ with a differentiable $F_{\varepsilon}^{\lambda}(x)$ for all $x\in X$. The following sections will achieve this task. In Section III.2, we show that $F_{\varepsilon}^{\lambda}(x)$ is the convex conjugate of an energy functional $E_{\varepsilon}^{\lambda}(v)$ that includes the field energy of $v$ (if appropriately defined), while Section III.3 identifies every $x\in X$ as a mix of internal and external densities. In the section we will only require Assumption 1.

### v2

@@ II.4 Formal density-functional theory
But it is also known since a long time that $v$-representability, even if it is secured for a dense set, is a problematic notion since most densities $\rho\in D$ will not have this property [30; 103]. In many cases (including standard DFT), the set of $v$-representable densities is known only implicitly (exceptions are lattice systems [25; 79; 80] and particles on a one-dimensional torus [94; 24]). Even worse, in the standard setting of DFT with density space $L^{1}\cap L^{3}$, arbitrarily close to each $v$-representable density there lies a non-$v$-representable one, which makes $F^{\lambda}(\rho)$ discontinuous and non-differentiable [59]. Such an irregular behavior of the universal functional $F^{\lambda}(\rho)$ will create complicacies in the optimization procedures that DFT relies on. Further, full (interacting and non-interacting) $v$-representability is imperative for the construction of the KS iteration scheme (see Section II.5). If, on the other hand, $F^{\lambda}(\rho)$ would be differentiable, the derivative $v=-\nabla F^{\lambda}(\rho)$ would yield an immediate solution to the $v$-representability question, because this potential $v$ then exactly gives $\rho$ as a ground-state density by Eq. (38). We will see in Section III how MY regularization offers a way out of this dilemma.
@@
The Hohenberg–Kohn theorem [47] states that, for a $v$-representable density $\rho\in D$, the corresponding potential is unique up to the addition of a constant. In standard DFT, a rigorous proof of the Hohenberg–Kohn theorem [35] has been achieved for $v\in L^{p}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$, with $p>2$, and it extends to closely related settings [36]. This leaves $3/2\leq p\leq 2$ as undecided (the lower bound is given by the argument for self-adjoint Hamiltonians from before), although results on unique continuation [48; 92] suggest that the Hohenberg–Kohn theorem holds also in this range. In non-standard DFT settings, like finite lattices or when currents are included, counterexamples are known [79; 78].
@@
Combining the assumptions given in this section and in Section II.1 (which will later serve as the basis for MY regularized DFT), we arrive at the following important result:
@@
Proposition. Assumption 1 together with Assumption 4 with a continuously embedded potential space $X^{*}\subset L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$ imply that every $v\in X^{*}$ supports a ground state—that is, for each $v\in X^{*}$, the superdifferential $\partial E^{\lambda}(v)$ is non-empty.
@@
This claim can be seen as follows. The choice $L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$ as a potential space makes $E^{\lambda}(v)$ continuous [63, Th. 3.1(iii)], implying that each continuously embedded $X^{*}\subset L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$ gives a continuous energy functional $E^{\lambda}(v)$ as well. But a continuous and concave functional on $X^{*}$ admits a non-empty subdifferential belonging to $X^{\ast\ast}$ everywhere in the interior of its domain (here all $X^{\ast}$) [8, Prop. 2.36]. In general, $X^{**}\supseteq X$, but with Assumption 1 we have reflexivity and thus $X=X^{**}$. By the additional requirement that $F_{\mathrm{DM}}^{\lambda}(\rho)$ is lower semicontinuous, we have $F^{\lambda}(\rho)=F_{\mathrm{DM}}^{\lambda}(\rho)$. Each $\rho\in\partial E^{\lambda}(v)$ is now in $D\subset X$ and through $v\in-\partial F^{\lambda}(\rho)=-\partial F_{\mathrm{DM}}^{\lambda}(\rho)$ it is connected to an actual ground-state solution of the Schrödinger equation as argued above. This demonstrates the claim of the proposition.
@@
We illustrate this subtle and important point with an example. Take the Hamiltonian from Eq. (23) with just a single particle and set $v_{\mathrm{off}}=v=0$. Then there is no bound state and thus $\partial E^{\lambda}(v)$ should be empty (note that the $\lambda$ parameter would not be needed for single particles since there are no interactions). Indeed, this system is incompatible with reflexivity and lower semicontinuity assumed together. If we give up on reflexivity and in contrast to Eq. (13) allow a superdifferential in $X^{**}$, then we can find an $x\in\partial E^{\lambda}(0)$ where $x\in X^{**}\setminus X$, which does not represent a ground state but instead corresponds to a scattering state. Next, we keep reflexivity but give up on lower semicontinuity of $F_{\mathrm{DM}}(\rho)$. We pick a minimizing sequence $\rho_{i}\to x$ in Eq. (30) consisting, for example, of normalized Gaussian-shaped densities $\rho_{i}\in D$ that get broader with increasing $i$, while $x\notin D$. Note that this means $x=0$, which cannot hold if $X$ includes the $L^{1}$ norm (and which would make it non-reflexive), since all $\rho_{i}$ were assumed normalized to the same number and clearly $N=\|\rho_{i}\|_{L^{1}}\to 0$ is a contradiction. But it is not necessarily a contradiction in a reflexive $X$ without the $L^{1}$ norm. Now the chosen sequence fulfills $\lim F_{\mathrm{DM}}(\rho_{i})=0$ but at the same time $F_{\mathrm{DM}}(x)=+\infty$, since $x\notin D$, so lower semicontinuity indeed cannot hold.
If, on the other hand, we add a fixed offset potential to the Hamiltonian that always guarantees existence of a ground state, for example, a harmonic oscillator potential $v_{\mathrm{off}}(\bm{r})=|\bm{r}|^{2}$, then the sequence $\rho_{i}$ of Gaussian-shaped densities from before cannot minimize $F_{\mathrm{DM}}(\rho)$ any more. This is because the definition of $F_{\mathrm{DM}}(\rho)$ includes the offset potential by Eq. (36) and, as the densities get broader, the potential energy $\langle v_{\mathrm{off}},\rho_{i}\rangle$ diverges. Every potential that has $\lim_{|\bm{r}|\to\infty}v_{\mathrm{off}}(\bm{r})=\infty$ guarantees existence of a ground state [84, XIII.47] and we call it a confining potential.
@@
This section summarized the most important features of DFT in the language of functional and convex analysis. The reader may wonder if Assumption 1 (or one of the two stronger ones) and Assumption 4 can be satisfied simultaneously. After all, the classical setting of DFT, with $X=L^{1}(\mathbb{R}^{3})\cap L^{3}(\mathbb{R}^{3})$ for densities and its dual $X^{\ast}=L^{3/2}(\mathbb{R}^{3})+L^{\infty}(\mathbb{R}^{3})$ for potentials, allows for Assumption 4 but breaks reflexivity, a property not needed for a consistent formulation of DFT but necessary for our later results related to Moreau–Yosida regularization. One solution is to choose a bounded domain $\Omega$, which implies a continuous embedding $L^{p}(\Omega)\subseteq L^{q}(\Omega)$, $q<p$. Given that the important result of lower semicontinuity due to Lieb [63, Th. 4.4 and Cor. 4.5] holds with respect to the $L^{1}$ topology, it extends to each $L^{p}(\Omega)$, $p>1$. We can then choose any $X=L^{p}(\Omega)$ with $1<p<\infty$ with dual $X^{*}=L^{p^{*}}(\Omega)$, $1/p+1/p^{*}=1$, and thus have Assumptions 2 and 4 simultaneously satisfied. By selecting $p=2$, we satisfy both Assumptions 3 and 4. The previous works of Eschrig [31] and Kvaal et al. [54] used this route, while many other works on the topic had to give up on at least one of the assumptions. Rather than switching to a bounded domain, we can use the offset potential $v_{\mathrm{off}}$ to introduce a confining potential for the whole system and thus ensure that every $v\in X^{*}$ yields a ground-state density. This, at the same time, serves to break the full translational and rotational symmetry of the setting, which is necessary to allow for any localized ground state in the first place [93].
Summary. To allow for any of Assumptions 13 together with 4 simultaneously, possible strategies are to switch to a bounded domain $\Omega\subset\mathbb{R}^{3}$ or to add a confining offset potential $v_{\mathrm{off}}$.
@@ II.5 Kohn–Sham method
DFT reformulates the ground-state problem of quantum mechanics by replacing the many-particle wave function or density matrix as the central variable with the one-particle density, thereby significantly reducing the complexity of the description. Yet, the usefulness of such a reduced description entirely depends on whether $F^{\lambda}(\rho)$ can be evaluated accurately and efficiently. The $F^{0}(\rho)$ of a non-interacting system amounts to just the kinetic energy and its conjugate functional $E^{0}(v)$ can be computed effectively by means of orbitals. For the interacting case, however, $F^{1}(\rho)$ is in principle unavailable since its evaluation corresponds to a quantum version of an NP-hard problem [90]. The difference $F^{1}(\rho)-F^{0}(\rho)$, which relates to the exchange–correlation functional, has moreover turned out to be difficult to model in terms of $\rho$ [100]. Thus, although the exact DFT formulation is very appealing by avoiding the full description by a wave function, further steps are needed to attain a useful reformulation. In the KS method [50], an auxiliary system of non-interacting electrons is introduced and its external potential $v_{\mathrm{s}}$ is chosen in such a way that it reproduces the same density as the interacting system. Yet, we will argue below that it is by no means clear that this approach is always possible.
The exchange–correlation potential is defined as the potential that must be added to the external potential and the Hartree potential of the auxiliary, non-interacting system in order to get the same ground-state density as the interacting system. This means that the method requires $\rho\in D$ to be simultaneously interacting and non-interacting $v$-representable, which is equivalent to $\partial F^{1}(\rho)$ and $\partial F^{0}(\rho)$ both non-empty. We already noted on the general issues regarding $v$-representability in Section II.4, and this additional requirement of simultaneous $v$-representability comes as a further restriction [101]. Assuming for now that this is ensured, we can take $v\in-\partial F^{1}(\rho)$ and $v_{\mathrm{s}}\in-\partial F^{0}(\rho)$ and define the exchange–correlation potential
@@
which itself is the functional derivative of the Hartree mean-field energy
@@
$v_{\mathrm{off}}+v_{\mathrm{s}}=v_{\mathrm{off}}+v+v_{\mathrm{H}}(\rho)+v_{\mathrm{xc}}(\rho),$ (43)
@@
The second step corresponds to solving the non-interacting Schrödinger equation with an external potential $v_{i+1}$ and determining a ground-state density $\rho_{i+1}$. If the Hamiltonian is defined on a bounded domain $\Omega\subset\mathbb{R}^{3}$ or contains a confining offset potential $v_{\mathrm{off}}$, this approach is always possible. A different convex-analytical argument for the non-emptiness of the superdifferential is given by our proposition building on Assumption 1 together with Assumption 4 and this was similarly used before [58, proof of Th. 12]. The two steps of Eqs. (44) and (45) can be combined into a single step, either for the potentials or, as we will do here, for the densities,
@@
For initialization, we choose an initial density $\rho_{1}$ or, equivalently, an initial potential $v_{1}$ for the auxiliary system that roughly leads to the desired ground-state density $\rho$. Since the auxiliary system does not include any interactions between the particles, we choose (in the simplest case) its ground state as a Slater determinant and solve Eq. (45) by filling up the orbitals of a one-particle Schrödinger equation with external potential $v_{\mathrm{off}}+v_{i+1}$. If the procedure converges, we get
@@
so indeed both systems share the same ground-state density if, in the auxiliary system, an external potential $v_{\mathrm{off}}+v_{\mathrm{s}}$ is applied. Finally, in order to conveniently retrieve the interacting energy $E^{1}(v)$ from a converged density and the expressions defined above, we can just use $E^{0}(v_{\mathrm{s}})=F^{0}(\rho)+\langle v_{\mathrm{s}},\rho\rangle$ and calculate
@@
where we have introduced the exchange–correlation functional $E_{\mathrm{xc}}(\rho)=F^{1}(\rho)-F^{0}(\rho)-E_{\rm H}(\rho)$. We remark that practical KS-DFT methods are usually based on approximating $E_{\mathrm{xc}}(\rho)$ and assuming differentiability, we then have $v_{\mathrm{xc}}=\nabla E_{\mathrm{xc}}(\rho)$.
Note that this formulation of the KS method does not come with any guarantee of convergence. Indeed, in practice, an additional damping step [49; 22; 23], like taking $\rho_{i}+\alpha(\rho_{i+1}-\rho_{i})$ with $\alpha<1$ as the next density, is introduced to facilitate convergence. A mathematical proof of convergence has only been achieved in the MY-regularized setting, which is in any case needed to avoid the $v$-representability problem, as will be discussed in Section V.3.
@@ III Moreau–Yosida regularization in density-functional theory
Having the stage set, we now present three different ways to introduce MY regularization into DFT, which are equivalent and thus have exactly the same outcome—namely, replacing the universal density functional $F^{\lambda}(\rho)$ with a differentiable $F_{\varepsilon}^{\lambda}(x)$ for all $x\in X$. Which way is to be preferred, is up to the reader, but we think it is illustrative to put them all side to side. Some readers might find one or the other easier to understand or more or less conceptually appealing. In Section III.2, we show that $F_{\varepsilon}^{\lambda}(x)$ is the convex conjugate of an energy functional $E_{\varepsilon}^{\lambda}(v)$ that includes the field energy of $v$ (if appropriately defined), while Section III.3 identifies every $x\in X$ as a mix of internal and external densities. In this section, we only require Assumption 1, but it is useful to also have Assumption 4 with $F^{\lambda}=F_{\mathrm{DM}}^{\lambda}$ fulfilled, so that no ambiguity about the respectively used universal functional arises.